OpenStudy (anonymous):

x^2 + ax + b = (x + c) (x + d) how do you get c and d from a and b?

OpenStudy (tkhunny):

Why not go the other way? b = c*d a = ???

OpenStudy (anonymous):

I'm trying to understand factoring and we've been starting fro x^2 + ax + b = 0

OpenStudy (tkhunny):

Well, it does not hurt to work backwards. Very often, this leads to greater understanding. (x + c) (x + d) = x^2 + (c+d)x + cd Thus a = c+d and b = c*d. These are very important relationships. You must find FACTORS of b that will SUM to a. This is the heart of factoring.

OpenStudy (anonymous):

thank you! exactly what I was looking for. that really helps.